In Drosophila, a cross between triply abnormal females with the X-linked recessive traits yellow bodies (y), echinus bristles (ec), and white eyes (w) and phenotypically normal males produces phenotypically normal F1 females. These F1 females are then testcrossed, and 1000 offspring occur as follows:

Body

Bristles

Eyes

number

normal

normal

normal

447

yellow

echinus

white

442

yellow

normal

normal

18

normal

echinus

white

17

yellow

normal

white

28

normal

echinus

normal

33

normal

normal

white

8

yellow

echinus

normal

7

What is the order of these three loci on the X chromosome?                            [ Select ]                       ["y, ec, w", "ec, y, w", "y, w, ec", "None of the answer options is correct."]      

What is the distance between the genes encoding body color and eye color?                            [ Select ]                       ["1.5 cM", "5.0 cM", "6.1 cM", "3.5 cM", "7.6 cM", "9.6 cM", "11.1 cM"]      
Answer #1

Answer:

The order of these three loci on the X chromosome = y, w, ec

The distance between the genes encoding body color and eye color = 5.0 cM

Explanation:

Body

Bristles

Eyes

number

Genotype

normal

normal

Normal

447

+++

yellow

echinus

White

442

y ec w

yellow

normal

Normal

18

y + +

normal

echinus

White

17

+ ec w

yellow

normal

White

28

y + w

normal

echinus

Normal

33

+ ec +

normal

normal

White

8

+ + w

yellow

echinus

Normal

7

y ec +

Hint: Non-recombinant (parental) combinations are produced more in number than recombinants. The parental (non-recombinant) genotypes is + + + / y ec w

1).

If single crossover occurs between y & ec…

Normal combination: + + / y ec

After crossover: + ec / y +

+ ec progeny=17+33 = 50

y + progeny = 28+18 = 46

Total this progeny = 96

Total progeny = 1000

The recombination frequency between y&ec = (number of recombinants/Total progeny) 100

RF = (96/1000)100 = 9.6%

2).

If single crossover occurs between ec & w

Normal combination: + + / ec w

After crossover: + w / ec +

+ w progeny= 28+8 = 36

ec + progeny = 33 + 7 = 40

Total this progeny =76

The recombination frequency between ec & w = (number of recombinants/Total progeny) 100

RF = (76/1000)100 = 7.6%

3).

If single crossover occurs between y & w…

Normal combination: + + / y w

After crossover: + w / y +

+ w progeny=17+8 = 25

y + progeny = 8+18 = 26

Total this progeny = 51

Total progeny = 1000

The recombination frequency between y&w = (number of recombinants/Total progeny) 100

RF = (51/1000)100 = 5.1%

Recombination frequency (%) = Distance between the genes (cM)

y--------5.1cM------w----7.6cM----ec

The order of genes = yellow----white---echinus